Larkspur Inc. is building a new hockey arena at a cost of $3,050,000. It received a down payment of $610,000 from local businesses to support the project, and now needs to borrow $2,440,000 to complete the project. It therefore decides to issue $2,440,000 of 10-year, 10.5% bonds. These bonds were issued on January 1, 2020, and pay interest annually on each January 1. The bonds yield 10% to the investor and have an effective interest rate to the issuer of 10.4053%. (There is an increased effective interest rate due to the capitalization of the bond issue costs.) Any additional funds that are needed to complete the project will be obtained from local businesses. Larkspur Inc. paid and capitalized $61,000 in bond issuance costs related to the bond issue. Larkspur prepares financial statements in accordance with IFRS. (USING PV TABLES)

a) Using (1) factor tables, (2) a financial calculator, or (3) Excel function PV, calculate the value of the bonds and prepare the journal entry to record the issuance of the bonds on January 1, 2020. (Hint: Refer to Chapter 3 for tips on calculating. For the journal entry, use the amount arrived at using the time value of money tables.)

b) Prepare a bond amortization schedule up to and including January 1, 2025, using the effective interest method.

c) Assume that on July 1, 2023, the company retires half of the bonds at a cost of $1,299,000 plus accrued interest. Prepare the journal entries to record this retirement.
Answer #1

Requirement a: Compute the value of bonds as follows

Particulars Amount
Present value of principal ($2,440,000 × 0.38554) $940,718
Present value of interest ($256,200 × 6.14457) $1,574,239
Selling price of the bonds $2,514,957

Note: Present value factor is 0.38554 when rate = 10% and number of periods = 10.

Present value annuity factor is 6.14457 when rate = 10% and number of periods = 10

Interest is $256,200 ($2,440,000 × 10.5%).

Prepare the following journal entry to record the bonds issued

Date Account Title and Explanation Debit Credit
Jan 1 Cash ($2,514,957 − $61,000) $2,453,957
2020 10.5% Bonds Payable $2,453,957
To record 10.5% bonds issued at premium

Note: Deduct bond issue cost of $61,000 from the proceeds of issue.

Requirement b: Prepare the bond amortization schedule as follows

Date

Cash Payment

(a)

Interest Expense

(b)

Discount Amortization

(a)(b)

Carrying Amount
Jan. 1 2020 $2,453,957
Jan. 1 2021 $256,200 ($2,453,957 × 10.4053%) = $255,342 $858 $2,453,099
Jan. 1 2022 $256,200 ($2,453,099 × 10.4053%) = $255,252 $948 $2,452,151
Jan. 1 2023 $256,200 ($2,452,151 × 10.4053%) = $255,154 $1,046 $2,451,105
Jan. 1 2024 $256,200 ($2,451,105 × 10.4053%) = $255,045 $1,155 $2,449,950
Jan. 1 2025 $256,200 ($2,449,950 × 10.4053%) = $254,925 $1,275 $2,448,675

Note: $2,453,099 ($2,453,957 − $858) ; $2,452,151 ($2,453,099 − $948) and so on.

Requirement c: Prepare the following journal entries to record the retirement of bonds

Journal entry for accrued interest

Date Account Title and Explanation Debit Credit
Jul 1 Interest Expense ($64,050 - $289) $63,761
2023 Bonds Payable ($1,155 × 1/2 × 1/2) $289
                 Cash ($256,200 × 1/2 × 1/2) $64,050
To record accrued interest on retired bonds

Journal entry for the reacquisition of bonds

Date Account Title and Explanation Debit Credit
Jul 1 Bonds payable ($2,440,000 ÷ 2) $1,220,000
2023 Loss on redemption of bonds $73,736
Bonds payable ($2,450,527 − $2,440,000) ÷ 2 $5,264
                   Cash $1,299,000
To record reacquisition of bonds

Notes: Compute carrying amount of bonds on July 1, 2013 as follows

Particulars Amount
Carrying amount of bonds - January 1, 2023 $2,451,105
Deduct: Amortization of bond premium ($1,155 ÷ 2) ($578)
Carrying amount of bonds - July 1, 2023 $2,450,527

Compute loss on redemption as follows

Particulars Amount
Reacquisition price of bonds $1,299,000
Deduct: Carrying amount of bonds - July 1 2023 ($2,450,527 ÷ 2) ($1,225,264)
Loss on redemption of bonds $73,736